OpenStudy (anonymous):

2013 Related Questions

OpenStudy (anonymous):

type any question related to the number 2013 then we will solve it together

OpenStudy (anonymous):

\[\int\limits_{}^{}e^{x^{2013}} dx\]lol jk

OpenStudy (anonymous):

\[\fbox{1} \]What is the remainder when \[1^2+2^2+...+2013^2\] is divided by 7

OpenStudy (anonymous):

lol @Euler271 i meant solveable questions,lol that one will kill me cos its not elemtntary function

OpenStudy (kc_kennylau):

@Jonask xD

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

for now dont post anything before the current question is solved,thank you

OpenStudy (kc_kennylau):

\[1^2+2^2+\cdots+2013^2\]\[=\frac16\times2013\times2014\times4027\]\[=671\times1007\times4027\]\[\equiv(-1)\times(-1)\times2\hspace{50pt}(\mbox{mod }7)\]\[=2\]

OpenStudy (anonymous):

ok so we used\[\sum i^2=\frac{n(n+1)(2n+1)}{6}\] awesome thanks @kc_kennylau

OpenStudy (kc_kennylau):

no problem :)

OpenStudy (anonymous):

For positive integers n, let the numbers \[c(n)\] be determined by the rules \[c(1)=1,c(2n)=c(n)\] and \[c(2n+1)=(-1)^nc(n)\]. Find the value of \[ \sum_{n=1}^{2013}c(n)c(n+2). \]

OpenStudy (anonymous):

basically it means for even numbers we have a number then for odd we have its negetive,mening most stuff cancel?

OpenStudy (kc_kennylau):

\[c(1)=1\]\[c(2)=c(1)=1\]\[c(3)=-c(1)=-1\]\[c(4)=c(2)=1\]\[c(5)=c(2)=1 \]\[c(6)=c(3)=-1\]\[c(7)=-c(3)=1\]lemme think about this later

OpenStudy (kainui):

http://math.unt.edu/problem-of-the-month scroll down to October 2013.

OpenStudy (anonymous):

\[\fbox{3}\]A positive integer n is such that n(n+2013) is a perfect square.

OpenStudy (anonymous):

a) Show that n cannot be prime.

OpenStudy (kc_kennylau):

lolz i don't know how to do i'm so dumb

OpenStudy (raden):

n(n+2013) = k^2 n^2 + 2013n - k^2 = 0 by using quadratic formula, we get n : n = (-2013 + sqrt(2013^2 - 4(1)(-k^2))) /2 n = (-2013 + sqrt(2013^2 + 4k^2)) /2 see n must be positive integer, therefore 4k^2 + 2013^2 = m^2 m^2 - (2k)^2 = 2013^2 (m+2k)(m-2k) = (3 * 11 * 61)^2 so far so good ? :)

OpenStudy (raden):

this way i get from you, @mukushla :)